OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

@dan815 @Nnesha @deepika.comet Please help!

OpenStudy (anonymous):

To be 100% correct, part A must say and non zero multiple

OpenStudy (anonymous):

@TuringTest Please help!

OpenStudy (anonymous):

@tanyajoseph99 proceed with equation one, multiply any non zero number with it. then re write the equation

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

-1(-3x + 7y )= (-16)*-1 =>3x-7y=16 --(1) -9x + 5y = 16---(2) (1) & (2) are our new system

OpenStudy (anonymous):

like this you can proceed just follow the instruction.! what is your difficulty on that ?

OpenStudy (anonymous):

-3x+7y=-16----> (1) -9x+5y=16 -----> (2) add 1 and 2 ==> -12x+12y=0 ==> -x+y=0 ==>x=y